OpenStudy (anonymous):

How do I determine sin theta to 3 decimal places when tan=1/2 and theta is in the first quadrant?

OpenStudy (anonymous):

Can you not provide the solution? just method to getting it?

OpenStudy (anonymous):

Thank you:)

OpenStudy (experimentx):

tan A = p/b p = 1, b = 2 sin A = p / (p^2+b^2)^(1/2)

OpenStudy (experimentx):

1/sqrt(5)

OpenStudy (anonymous):

How would I go about solving this qustion?

hero (hero):

Hint \[\tan \theta = \frac{opp}{adj}\] 

OpenStudy (anonymous):

How can I solve this?

OpenStudy (experimentx):

1/cotA = tan A = p/b find p and b from sinA read theta for A